OpenStudy (anonymous):

What is the value of the expression 8–2? A. –16 B. 1/64 C. 1/16 D. 64

jimthompson5910 (jim_thompson5910):

I'm assuming you meant to say \[\Large 8^{-2}\]

jimthompson5910 (jim_thompson5910):

If so, then use the rule \[\Large x^{-k} = \frac{1}{x^{k}}\]

OpenStudy (anonymous):

Yeah I'm sorry i just copied and pasted

OpenStudy (anonymous):

64?

jimthompson5910 (jim_thompson5910):

it's ok

OpenStudy (anonymous):

64?

jimthompson5910 (jim_thompson5910):

8^2 = 64, but \(\Large 8^{-2}\) is not 64

jimthompson5910 (jim_thompson5910):

\[\Large 8^{-2} = \frac{1}{8^{2}} = ???\]